OpenStudy (anonymous):

Type the equation of the given line in standard form. The line including two points (6, 4) and (-5, 3). Thanks for the help!

OpenStudy (anonymous):

Line with two points is given by: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x- x_1}{x_2 - x_1}\] Use this to solve...

OpenStudy (anonymous):

But I only have two y points. But I have to substitute in 3? Sorry, I'm just confused..

OpenStudy (kaiz122):

these are the points in your question \[ (x _{1},y_{1}) (x _{2},y_{2})\] just substitute to the formula

OpenStudy (anonymous):

y-4=((3-4)/(-5-6))(x-6) y-4=(1/11)(x-6)

OpenStudy (anonymous):

the equation of the line passing through two points (x1,y1) and (x2,y2) and slope m is: y-y1=m(x-x1) where m is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

@kaiz122 I know to substitute those in. But in the topmost equation, there is just plain x & y and I don't know what to put into those. Does that make sense?

OpenStudy (kaiz122):

those are constants @theequestrian

OpenStudy (kaiz122):

and nitz have already given the answer

OpenStudy (anonymous):

Ohhhh. Right. That was quite embarassing. Lol. Thanks though :)

OpenStudy (kaiz122):

np. :)

OpenStudy (anonymous):

Leave x and y as such they are arbitrary points.

OpenStudy (anonymous):

\[\frac{y-4}{3 - 4} = \frac{x-6}{-5-6} \implies \frac{y-4}{-1} = \frac{x-6}{-11}\] Can you solve it now??

OpenStudy (anonymous):

Oh, okay. & I think I can. :)

OpenStudy (anonymous):

Go ahead..

OpenStudy (anonymous):

I got y-4=(1/11)(x-6) once I solved it, like what nitz said.

OpenStudy (anonymous):

multiply by 11 both the sides and tell me what do you get..

OpenStudy (anonymous):

I got 11y-44=x-6

OpenStudy (anonymous):

No, wait. I got 11y-44=11x-66

OpenStudy (anonymous):

Previous one is right..

OpenStudy (anonymous):

On right hand side 11 and 11 will get cancel so need to multiply 11 with it also.. Getting??

OpenStudy (anonymous):

But don't I have to distribute the 11 to the x-6 too?

OpenStudy (anonymous):

Why you distribute ?? there is 11 in the denominator .. Don't you think both the 11 will get cancel??

OpenStudy (anonymous):

Getting ?? tell me..

OpenStudy (anonymous):

Oh. I thought it would only cancel for the 1/11 * 11. But I see what you are saying now

OpenStudy (anonymous):

Or I will show you??